OpenStudy (anonymous):

Find the simplified, exact value of each expression. cos(-30) sin(-pi/2) sin(-13pi/6)

Nnesha (nnesha):

negative angle identities \[\rm \sin(-x) = -\sin(x) ~~~\color{red}{odd}\]\[\rm \cos(-x)= \cos(x)~~~~~\color{red}{even}\]\[\rm Tan(-x)=-\tan(x)~~\color{red}{odd}\]

Nnesha (nnesha):

and then you can use unit circle to find EXACT value of each ratio or 30-60-90 theorem

Nnesha (nnesha):

for the last one first you should apply the negative identity rules and then subtract 2pi( 2pi is one complete circle ) so subtract 2pi from 13pi/2 simplify and then use the unit circle to find EXACT value